### v2

# Scalable Ensembling For Mitigating Reward Overoptimisation

Abstract
Reinforcement Learning from Human Feedback (RLHF) has enabled significant advancements within language modeling for powerful, instruction-following models. However, the alignment of these models remains a pressing challenge as the policy tends to overfit the learned “proxy” reward model past an inflection point of utility as measured by a “gold” reward model that is more performant – a phenomenon known as over-optimization. Prior work has mitigated this issue by computing a pessimistic statistic over an ensemble of reward models, which is common in Offline Reinforcement Learning but incredibly costly for language models with high memory requirements, making such approaches infeasible for sufficiently large models. To this end, we propose using a shared encoder but separate linear heads. We find this leads to similar performance as the full ensemble while allowing tremendous savings in memory and time required for training for models of similar size.
1 Introduction
Modern language models have been ubiquitous in discussions of general-purpose AI systems that can accomplish myriad tasks across many disciplines and with rapidly increasing capabilities OpenAI (2023); Manyika & Hsiao (2023); Touvron et al. (2023); Bommasani et al. (2022); Wei et al. (2022). However, alongside this increase in capabilities, there has been growing concern around the risks of such systems as they are not entirely interpretable and could be misused to cause substantial harm either maliciously or inadvertently (Hendrycks et al., 2023; Wang et al., 2023; Hendrycks et al., 2022).
A salient research question is that of alignment; given a set of human values, how do we imbue these principles within the behavior of such systems? Russell (2022). Suppose we can access a ground truth reward model. One observed phenomenon is that training a reward model from preference feedback eventually reaches an inflection point where increasing the policy performance on the proxy stagnates and ultimately degrades the reward credited by the ground truth model – even though both are consistent with human labels! (Gao et al., 2023).
This reward over-optimization problem has been identified as a significant technical issue in scaling up learning from human feedback Ouyang et al. (2022); Casper et al. (2023); Coste et al. (2023); Eisenstein et al. (2023). Prior work has noted that because the reward models trained from user feedback are only a proxy for their underlying preferences, opting for a high reward can lead to performance degradation. Recent work has proposed mitigations either through ensembling or transforming the RL objective. Still, these methods are either too computationally expensive or analytically intractable to be used with modern language models given limited compute (Coste et al., 2023; Eisenstein et al., 2023).
In our work, we propose a simple modification to a common strategy of ensembling: instead of maintaining multiple separate reward models, we opt for a shared backbone with different linear heads with the hypothesis that different initialization and training procedures will generate enough diversity. Our contributions show initial experiments validating the utility of such an approach in RLHF, as it is just as performant as using a full ensemble for mitigating overoptimization while requiring less time for training during reward modeling and less memory and time for PPO.
2 Related Work
Reinforcement Learning from Human Feedback (RLHF) RL is a powerful framework for learning diverse and performant policies that can tackle a wide array of tasks Sutton & Barto (2018); Abbeel & Ng (2004); Bellman (1957); Haarnoja et al. (2017); Drake (2005). Of particular recent interest are language models, which have shown impressive base capabilities in terms of instruction following. RLHF is a core component of the tuning processes for many of the currently highest-performing and most widely-used language models, improving their ability to follow instructions and align their responses with human preferences Stiennon et al. (2022); Christiano et al. (2023). However, tuning through RLHF can introduce a risk of overfitting to a proxy of true reward since the reward model is learned. Prior work analyzes this phenomenon using a fixed “gold-standard” reward model in place of humans, training proxy reward models from labels it provides (Gao et al., 2023).
Ensembles for Overoptimisation Recent work has shown multiple ways to mitigate this overoptimization issue, but the suggestions only apply during training or require expensive copies of multiple reward models for ensembling without deeper analysis as to why they mitigate this issue Coste et al. (2023). Related work further discovered that the ensembles are more effective when pretraining from scratch with separate seeds Eisenstein et al. (2023). Separately, other work has investigated uncertainty across ensemble members for RLHF, showing that using the same backbone with different linear heads can lead to significant improvement in calibration but that this weakly correlated with performance for summarization tasks Gleave & Irving (2022). Concurrent work investigates a similar approach through uncertainty estimation to propose a last layer reward modeling scheme with Bayesian uncertainty estimation Zhang et al. (2024b). Alternatively we use a linear layer for each separate reward function in an ensemble and train the reward model by fine-tuning through the base encoder in the language model. The recent work most similar to our approach investigates using either LoRA for fine-tuning a frozen transformer or similary fine-tuning linear heads with the same transformer backbone for a reward modeling ensemble Zhang et al. (2024a). Our approach differs in that we forgo using the mean or a least confidence bound estimate due to the observation that the min is better suited for minimizing epistemic uncertainty and is simpler to implement. This allows for utilizing the shared features from supervised fine-tuning while enabling the diversity of reward ensembles in a scalable manner with a simple minimum function.
3 Background & Methods
3.1 PPO
Proximal Policy Optimization (PPO) is an algorithm in reinforcement learning (RL) that is particularly adept at ensuring gentle policy updates, crucial for the complex dynamics of language models by optimizing reward with a KL constraint to the original policy.
$\max_{\pi}\underset{s,a\sim\pi_{\text{old}}}{\mathbb{E}}\left[\frac{\pi(a|s)}{\pi_{\text{old}}(a|s)}R(s,a)-\beta\text{KL}(\pi_{\text{old}}||\pi)\right]$ (1)
3.2 Reward Learning
In standard reward modelings given a prompt completion $x$ and a human pairwise binary preference label $y$, we instantiate the reward model using the backbone feature extractor from a pre-trained or model $\mathcal{F}$. We then initialize a linear reward head $\mathcal{H}:\mathbb{R}^{d}\times\mathbb{R}$ with feature dim $d$ giving a reward function for each linear head.
$r_{i}(x)=\mathcal{H}_{i}(\mathcal{F}(x))$ (2)
3.2.1 Multi-head Reward Learning
In our approach, the multi-head reward model is structured upon a shared base neural architecture derived from the pre-trained and supervised fine-tuned language model. Everything is fixed except instead of a singular head we design the model to incorporate multiple heads. Formally let $\mathcal{F}$ is the feature extractor and $\mathcal{H}_{i}$ is the $i^{th}$ head linear head, the reward $r_{i}$ for input $x$ using head $i$ is simply the minimum reward from the linear heads:
$\hat{r}(x)=\min_{i}\mathcal{H}_{i}(\mathcal{F}(x))$ (3)
Figure 1: Comparison of Reward Modeling methods to ours (right)
4 Experiments
4.1 Datasets and Motivation
We select the Alpaca Instructions dataset, which emphasizes naturalistic, closed-form questions and answers, usually having a well-defined correct response, e.g., the number of planets in the solar system, but also included more open questions such as famous actors who started on broadway, why certain states are named, how to play kickball, etc. Taori et al. (2023). Given the most relevant prior work studying overoptimisation focused on this dataset, it is our primary dataset for training and evaluation (Coste et al., 2023).
4.2 Methodology and Tools
We follow the RLHF pipeline of supervised fine-tuning (SFT), reward learning, and proximal policy optimization (PPO). The experiments use the OPT model family Zhang et al. (2022). We opted for the Alpaca Farm codebase as our framework (Dubois et al., 2023).
  • Supervised Fine-Tuning (SFT): Given a prompt (”Tell me a bedtime story”) and some ideal completions, the base model is fine-tuned to minimize perplexity on a split of 52k instructions.
  • Reward Learning: Given the SFT model and the same set of prompts but now with pairs of completions (preferred and dispreferred), we use the backbone and fine-tune a linear head on the Bradley-terry loss with the preferred completion as the target (Bradley & Terry, 1952).
  • PPO Finally, we further tune the SFT language model as a policy in the RL framework against the reward model through PPO.
We modify this pipeline by training the multi-head reward models with the base model initialized from SFT. When selecting the reward for a given sample, since we produce multiple predictions, we are presented with multiple choices for computing our final prediction Coste et al. (2023), but we simply take the minimum over the ensemble. A pessimistic estimator might reduce performance in principle. However, for offline RL and RLHF preferences, it has been either conjectured or empirically validated that a minimum is optimal, and prior work demonstrated uncertainty penalties do not improve the efficiency of ensembles in preventing over-optimization over a simple min Zhu et al. (2023); Coste et al. (2023); Kumar et al. (2020); Chen et al. (2021). Our approach is shown in Figure 1, and all hyperparameters and further training details are in the Appendix.
Figure 2: Gold analysis on top, Proxy metrics below.
4.3 Results
To test the efficacy of our approach, we run 1) Standard PPO 2) PPO with an ensemble of three reward models and 3) PPO with a multi-head reward model with three linear heads. We fix ensemble size at three as prior work found no gains when increasing to four or five members due to fixed compute, although we emphasize our method is amenable to a much larger number of ensemble members with minimal overhead Coste et al. (2023). We select the 1.3 B parameter OPT model as the proxy and the 6.7B variant as the gold model and run PPO for 15 epochs to extensively test overoptimisation. We also use the 1.3B model as the policy after supervised fine-tuning.
We present our results in Figure 6. Given the difference in reward scales and multiple works validating the benefit of using a full ensemble for overoptimisation, we opt for two figures that demonstrate 1) the efficacy of the multi-head ensemble against standard PPO and 2) the efficacy of using a multi-head or full ensemble. Following prior work, we plot reward against KL divergence Gao et al. (2023); Coste et al. (2023); Eisenstein et al. (2023). Figure 6 clearly shows the replication of overoptimisation as we reproduce the concave-down nature of gold rewards under standard PPO and then summarily show how using a multi-head reward model bridges this gap. Figure 6 shows that the multi-head and full ensemble nature to similar gold performance however we emphasize that our approach allows an ensemble with much larger reward models and following the prior work suggesting the full ensemble we train each member for three epochs to maximize performance whereas we find the multi-head approach needs only one epoch. We average over three seeds for each method and include further details and ablations in the appendix.
4.4 Calibration
Figure 3: Multi-head model calibration over different objectives with respect to the probabilities (taking min, max over ensemble etc.)
4.5 Calibration Analysis and Implications
We investigate the calibration of our multi-head reward models, as an accurate representation of uncertainty is crucial in RLHF to prevent overoptimistic predictions Casper et al. (2023). Prior work found that ensembling improves calibration but is weakly related to model error when using a shared backbone with different linear heads, suggesting that separate reward models might be necessary for efficient ensembling Gleave & Irving (2022). However, we hypothesize that the difference in improvement is due to the diversity of tasks, as the AlpacaFarm dataset focuses on more open-ended questions compared to the summarization tasks in the prior work, which might increase model error due to epistemic uncertainty. We assess the calibration quality of the models using different ensemble objectives, such as mean and minimum reward estimation. The multi-head models, particularly with the minimum objective, demonstrate an enhanced ability to capture uncertainty in the reward signals without incurring significant computational overhead. While the minimum estimate is technically miscalibrated, it leads to better-than-predicted performance at high levels of certainty. This suggests a regularization effect from the pessimistic estimator, as overoptimized models typically reach a local minima, explaining the ”U-shaped” curves that result as KL divergence from the base policy increases. Our findings align with previous work on pessimistic estimates in ensemble models, underscoring the benefits of a conservative approach in certain RL contexts. Furthermore, our experiments suggest that a smaller number of heads (3) might be optimal for mitigating overoptimisation, providing a valuable guideline for future research in ensemble-based models for RLHF. An increase in the number of heads beyond a certain threshold might introduce more noise than beneficial diversity, potentially leading to overfitting on dataset nuances, which is consistent with existing literature on the effectiveness of smaller ensembles in specific scenarios Galdran et al. (2023). See Figure 3 for details.
4.6 Conclusion
Our research contributes a novel method to improve the robustness of aligning language models by utilizing a robust pessimistic statistics while avoiding the cost of materializing a full ensemble. By addressing the challenges of overoptimization computationally efficiently, we move a step closer to developing AI systems that can reliably align with human values and preferences. We plan to further investigate the effects of different datasets, such as the more naturalistic Stanford Human Preferences Shah et al. (2019) and larger model sizes for future work.
4.7 Acknowledgements
AMA acknowledges support from a Knight-Hennessy fellowship and a National Science Foundation fellowship. SK acknowledges support by NSF 2046795 and 2205329, NIFA award 2020-67021-32799, the Alfred P. Sloan Foundation, and Google Inc.
Appendix A Appendix
A.1 D Additional experimental details
A.1.1 D.1 Hyperparameters
We give the hyperparameters here for different components of our RLHF pipeline:
Table 1: SFT hyperparameters for 6.7B model
Parameter Value
Learning rate 2e-5
Epochs 3
Batch size 128
Table 2: SFT hyperparameters for 1.3B
Parameter Value
Learning rate 8e-6
Epochs 1
Batch size 128
Table 3: RM hyperparameters
Parameter Value
Learning rate 1e-5
Epochs 3
Batch size 64
Table 4: PPO hyperparameters
Parameter Value
Max instruction length 520
Max new tokens (answer length) 256
PPO epochs 4
Top-p 0.9 (1.0 for PPO training)
Top-k 0
Temperature 1.0
Rollout Batch size 512
Gradient Step Batch size 256
Learning Rate 6e-6
A.1.2 D.2 AlpacaFarm Dataset Details
The AlpacaFarm dataset Dubois et al. (2023) employed in our experiments uses the Alpaca data Taori et al. (2023) made up of 52,000 samples. This data is chosen due to its large size and success in training instruction-following models. AlpacaFarm contains five splits: a labeled 10k ”sft” split for supervised fine-tuning, a 10k ”pref” split containing pairwise preference labels, a 20k ”unlabeled” split for training algorithms such as PPO, a 2k validation split, and an unused 10k split. We use the noisy variant of alpaca instructions with 25% label noise to more closely model real-world data distributions.
A.1.3 Reward model training ablation
Interestingly, we find that if we only train each reward model in the full ensemble for one epoch the ensemble approach is not sufficient to prevent over-optimisation, and that at least three epochs are necessary while prior work had five or more. We hypothesize that the multi-head approach is more amenable to underspecified features for reward modeling.
Figure 4: Gold analysis.
Figure 5: Proxy metrics.
Figure 6: Gold analysis on top, Proxy metrics below.